OpenStudy (anonymous):

How can I write this as an algebraic expression, " Julia spent $10 on school supplies. This was two thirds of the total amount of money she earned for the week".

OpenStudy (texaschic101):

x = total amount earned for the week (2/3)x = 10 <== thats your algebraic expression (2/3)x = 10 -- divide both sides by 2/3 x = 10/(2/3) x = 10 * 3/2 x = 30/2 x = 15 check.. (2/3)x = 10 2/3(15) = 10 30/3 = 10 10 = 10 (correct) Just so you know, she earned $15 for the week and 2/3 of it ($10) was spent on school supplies.